OpenStudy (anonymous):

Derivative word problem: The motion of a spring that is subject to a frictional force or a dampening force (such as a shock absorber in a car) is often modeled by the product of an exponential function and a sine or cosine function. Suppose the equation of motion of a point on such a spring is s(t) = 3e^(-4.5t)sin (2πt), where s is measured in centimeters and t in seconds. Find the velocity after t seconds.

OpenStudy (anonymous):

derivative of fg = g'f + gf'. so if, f = 3e^(-4.5t) and g = sin(2πt) then... the derivative of f = 3e^(-4.5t) the derivative of g = cos(2πt) plug in respectively... s'(t) = (cos(2πt) * 3e^(-4.5t)) + (sin(2πt) * 3e^(-4.5t)) s'(t) = 3e^(-4.5t) * (cos(2πt) + sin(2πt)) The above should be your velocity, please correct me if I'm wrong.

OpenStudy (anonymous):

Great! Thanks!

OpenStudy (anonymous):

You're welcome